If 100 ml of hcl is neutralised by 100 ml 4M naoh the purity of concentrated hcl (specific gravity=1.2)

Grade:

1 Answers

Vikas TU
5 years ago
Dear Student,
as HCl and NaOH are the strongest acids and base respectively we can apply here:
M1V1 = M2V2
Therefore,
M*100 = 100*4
M = 4M
Hence purity is: (4/4)*100 = 100%